OpenStudy (anonymous):

^3√logx How to work out the derivative?

OpenStudy (anonymous):

\[\sqrt[3]{logx}\] This is the Math Notation.

OpenStudy (anonymous):

\[Let ...f(x)=\sqrt[3]{x}.....and ....g(x)=\log(x)\] Now can u tell me the derivative of f(g(x))?

OpenStudy (anonymous):

Hello...........

OpenStudy (anonymous):

Using the chain rule now to work this out.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1/3x^2/3(logx)1/x(x) ?

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }(\log x)^{\frac{ -2 }{ 3 }}{\times}\frac{ 1 }{ xln10 }\]

OpenStudy (anonymous):

Where did log10 come from?